### v1

# Enhanced premelting at the ice–rubber interface using all-atom molecular dynamics simulation

Abstract
The ice–rubber interface is critical in applications such as tires and shoe outsoles, yet its molecular tribology remains unclear. Using all-atom molecular dynamics simulations, we studied premelting layers at the basal face of ice in contact with styrene–butadiene rubber from 254 to 269 K. Despite its hydrophobicity, rubber enhances structural disorder of interfacial water, promoting premelting. In contrast, water mobility is suppressed by confinement from polymer chains, leading to glassy dynamics distinct from the ice–vapor interface. Near the melting point, rubber chains become more flexible and penetrate the premelting layer, forming a mixed rubber–water region that couples the dynamics of both components. These results suggest that nanoscale roughness and morphology of hydrophobic polymers disrupt ice hydrogen-bond networks, thereby enhancing premelting. Our findings provide molecular-level insight into ice slipperiness and inform the design of polymer materials with controlled ice adhesion and friction.
1 Introduction
The ice–rubber interface plays a key role in many industrial applications, including vehicle tires and shoe outsoles. For tires, rubber materials help reduce the risk of slipping and allow efficient transmission of engine power and braking force to icy road surfaces. Rubber friction arises from its viscoelasticity and surface adhesion on solid surfaces, typically showing a bell-shaped dependence on sliding velocity 1, 2, 3. In contrast, the tribology of rubber on ice differs markedly because of the unique properties of ice. Ice is exceptionally slippery, and its low friction has been attributed to several mechanisms, including frictional heating, premelting layers, pressure melting, and the formation of thin liquid-like surface films 4, 5, 6, 7, 8, 9. Nevertheless, the fundamental origin of this slipperiness remains unresolved.
A thin liquid-like layer, known as the premelting layer, forms on ice surfaces and strongly affects interfacial behavior 10, 11, 12. Its thickness has been reported from a few nanometers to micrometers, and although debated, it is widely accepted that the layer behaves like a liquid. Near the melting point, it is highly sensitive to temperature, where changes of only a few kelvins can strongly alter its morphology, structure, and dynamics 13, 14, 15, 16. Molecular simulations have been widely used to probe the structure and dynamics of premelting layers 17, 18, 14, 19, 20, 21, 16. Recent computational studies show that small molecules can locally alter ice surfaces. For example, ions disrupt the local ice structure 22, 23, and short hydrophilic polymers induce partial melting and suppress ice growth 24. However, the effects of polymer interfaces remain largely unexplored.
The ice–polymer (or rubber) interface has been studied experimentally for decades, demonstrating that the surface properties of ice are strongly influenced by the chemistry and microstructure of the contacting material. Friction measurements showed that the viscosity of the premelting film is sensitive to surface chemistry 25. Hydrophobic coatings reduce ice adhesion, whereas hydrophilic polymers interact with liquid-like films to form hydration layers 26, 27. Lecadre et al.28 suggested that friction is governed by the viscoelasticity of contacting materials at low temperatures, but dominated by premelting layers at higher temperatures. Molecular dynamics simulations further revealed that wettability strongly affects premelting-layer friction29, with similar behavior seen in supercooled water 30. More realistic interface models, including polymer and graphene interfaces, showed that hydrogen bonding with the contacting surface alters the orientation, structure, and dynamics of water molecules in the premelting layers 31. These findings suggest that both the chemical composition and nanostructure of rubber chains can modify premelting layers, producing behaviors distinct from those at the pure ice–vapor interface. However, the atomistic details of such rubber-induced effects remain largely unexplored.
Rubber is mainly composed of nonpolar polymers such as styrene and butadiene, often combined with cross-linkers and fillers. Its hydrophobic nature is expected to affect the ice surface through van der Waals interactions. In this study, we use molecular dynamics simulations to investigate the atomistic structure and dynamics at the ice–rubber interface, as slipperiness is closely linked to molecular mobility 32. We find that rubber enhances structural disorder in the premelting layers while suppressing their dynamics due to confinement. Importantly, we also show that the dynamics of interfacial rubber chains change cooperatively with premelting water molecules, and at near the melting temperature, a mixed rubber–water layer emerges.
2 Methods
2.1 Molecular Dynamics Simulation
[…]
Chemical species mol % wt %
Styrene 13.8 23.3
1,2-Butadiene 22.5 20.0
Cis-1,4 -butadiene 10.0 8.9
Trans-1,4 butadiene 53.7 47.8
Table 1: SBR composition
Figure 1: Simulation system. (A) Chemical structure of styrene-butadiene rubber used in this work. The number of monomer unit is indicated and sequence of monomer was randomly shuffled to . (B) Representative snapshot of the system. The top side is the ice–rubber interface, while the lower is the ice–vapor interface.
[…]
2.2 Density Profile
The density profile was computed based on atomic masses after re-centering the system on the inner bulk region of the ice crystal. A bin width of 0.02 nm was used for the analysis.
2.3 Diffusion coefficient
The molecular dynamics of premelting layers was evaluated using the mean square displacement (MSD),
$\begin{split}\mathrm{MSD}(\Delta)=\langle|\bm{r}(t+\Delta)-\bm{r}(t)|^{2}\rangle\end{split}$ (1)
where $\langle\cdot\rangle$ means the ensemble average, and $\bm{r}(t)$ is the position of the oxygen atom. In this calculation, we counted molecules that stayed in the combined region of layers 1 and 2 during the measurement time, $\Delta$. The diffusion coefficient, $D$, was computed from the sloop of MSD in the nearly linear region (after lag time 100 ps),
$D=\frac{1}{6}\frac{d}{dt}\mathrm{MSD}(t)$ (2)
2.4 Analysis of single molecular dynamics using machine learning
Although MSD is useful for characterizing system-averaged molecular dynamics, single-molecule dynamics cannot be fully captured. To distinguish liquid-like mobile molecules from solid-like immobile molecules, a machine-learning approach for single-molecular dynamics analysis has been developed 44, 45, 46. In this work, we employ this method to clarify the rubber-inducing effect on the premelting molecules at single-molecule resolution. In the machine learning approach, we first sample the short-term trajectories from the rubber-ice interface, $\bm{x}$, and those from bulk ice systems, $\bm{x^{\prime}}$. Any type of trajectories can be used for $\bm{x^{\prime}}$, but we employed a time series of the oxygen atom displacement in the water molecules for 64 ps, according to our previous work 46. Both $\bm{x}$ and $\bm{x^{\prime}}$ are input into a neural network, which outputs a score, $g(\bm{x})$, that quantifies the difference from the solid bulk water system,
$g(\bm{x})=\mathbb{E}_{\bm{x}^{\prime}\sim\bm{y}^{\prime}}\left[f^{*}(\bm{x})-f^{*}(\bm{x}^{\prime})\right]$ (3)
where $\mathbb{E}$ represents the expectation value taken over samples $\bm{x}$ sampled from the distribution of $\bm{y}$, and $f^{*}(\bm{x}^{\prime})$ is the function represented by the neural network, which is a simple multi-layer perceptron. The neural network model trained in ref. 46 was used. Higher values of $g(\bm{x})$ correspond to molecular dynamics that differ from the solid phase. Following our previous work, a threshold value of $g(\bm{x})$, which was determined based on the comparison of liquid and solid bulk systems, was used to distinguish liquid and solid molecules in the premelting layers.
2.5 Analysis of polymer dynamics
We quantified the interfacial fluctuations of rubber by calculating the root mean square fluctuation (RMSF) of non-hydrogen rubber atoms,
$\mathrm{RMSF}=\langle|\bm{r}(t)-\bar{\bm{r}}|\rangle,$ (4)
where $\bm{r}$ is the atom position, and $\bar{\bm{r}}$ is the time-averaged position of each atom. Rubber atoms within the range $79.05~\text{nm}\leq z\leq 80.06~\text{nm}$, which includes both the premelting–mixing layers and part of the bulk interfacial rubber, were used consistently across all temperatures. To evaluate short-term dynamics, the trajectory was segmented into 5 ns windows, and the RMSF was computed separately for each segment. Trajectories every 100 ps were used in this computation to remove highly correlated frames.
3 Results
3.1 Enhanced disorder of premelting layers at the ice–rubber interface
As premelting layers are highly sensitive to temperature near the melting point 13, 14, we compare the atomistic surface structures at 264 and 269 K for the ice–vapor and ice–rubber interfaces. Snapshots show that the surface structures are similar at 264 K, whereas at 269 K, the third layer becomes more disordered at the rubber interface, compared to the vapor interface (Fig. 2A). The density profiles show that rubber and premelting layers coexist in a very thin surface layer at both temperatures (Fig. 2B). However, the density profiles at 269 K show weaker peaks at the rubber interface compared to the vapor interface, while their profiles at 264 K are nearly identical (Fig. 2C). This difference is evident even in the 3rd ice layer, which is not directly exposed to rubber atoms. Interestingly, at 269 K, the third layer at the ice–rubber interface exhibits comparable peak heights and widths in the density profile to the second layer at the ice–vapor interface, indicating structural similarity between the two layers (Fig. 2C). Comparison of the profiles at different temperatures in the range of 254–269 K (Fig. S1) indicates a sharp increase in ice premelting between 267 and 269 K, accompanied by enhanced penetration of rubber atoms into the premelting layer.
Figure 2: Molecular structure of ice–rubber and ice–vapor interfaces. (A) Representative snapshots at 264 and 269 K for the ice–rubber and ice–vapor interfaces. Labels (1st, 2nd, and 3rd) indicate the layer positions (Layer 1–3). (B) Density profile of water oxygen atoms and rubber atoms in the ice–rubber interface at 264 and 269 K. (C) The density profile of water oxygen atoms, comparing the ice–rubber and ice–vapor interfaces, at 264 and 269 K.
Figure 3: Comparison of the dynamical properties of ice–rubber and ice–vapor interfaces. (A) Mean squared displacement (MSD) of premelting layers (combined Layer 1 and Layer 2) in the ice–rubber interface. $\alpha$ is the scaling exponent of MSD. (B) Diffusion coefficient as a function of temperature. (C) Distribution of the score, $g(\bm{x})$, which quantifies deviation from bulk solid behavior in single-molecule short-term dynamics. Layer 1–4 represent the distributions from the ice–rubber interface at 269 K, while ”Ice” and ”Water” correspond to bulk solid and liquid systems, respectively. The black dashed line indicates the threshold used to classify molecules as solid-like or liquid-like. (D–F) Ratio of liquid-like molecules in Layers 1–3 at different temperatures. Error bar shows standard deviation of the mean values computed using 5 ns time-blocks in (B) and (C–F).
3.2 Suppressed water dynamics by rubber confinement
Building on the observation of enhanced structural disorder at the ice–rubber interface, we use dynamical properties to clarify whether the premelting layers behave as mobile liquid. Given that the relaxation times of the first and second layers are on the order of several tens of picoseconds 18, 46, we treated these two layers as a single region and computed the MSD of molecules within the combined region. For the rubber interface, MSD increased significantly between 264 and 269 K (Fig. 3A). The scaling exponent, $\alpha$ ($\rm{MSD}\propto t^{\alpha}$), ranged in 0.85–0.89 for the vapor interface, and 0.77–0.87 for the rubber interface (Tab. S1), indicating the glassy behavior of the premelting layers 19. The ice–rubber interface showed a lower $\alpha$ value compared to the ice–vapor interface, suggesting that the presence of rubber molecules enhances the confinement of water molecules within the premelting layers.
To further clarify the temperature dependence of interfacial dynamics, we computed the diffusion coefficients for the ice–vapor and ice–rubber interfaces. As shown in Fig. 3B, the diffusion coefficient at the ice–vapor interface increases linearly with temperature, whereas the ice–rubber interface exhibits a change in the slope between 264 and 267 K. These results highlights distinct differences in both diffusivity and temperature dependence between the two interfaces.
Besides the mean behavior within the layer, we speculate that the local nanostructure of the rubber molecules contributes to the heterogeneity in the layer, thus requiring a single-molecular dynamics analysis. Here, from the measurement of the short-term molecular dynamics (64 ps) of ice premelting layers at the single-molecule resolution 46, we quantify the ratio of mobile-liquid-like molecules. We computed the score $g(\bm{x})$ for the four premelting layers, which represents differences from the bulk ice molecules on the short-term dynamics, and compared the score to those of bulk solid and liquid water systems. For the rubber interface, Layers 1 and 2 show similar properties to the bulk liquid system (Fig. 3C). The threshold to classify solid and liquid molecules was determined at 269 K (Fig. 3C) and subsequently applied to all other temperatures, following the approach used in previous work 46.
We compared the rubber and vapor interfaces in terms of the ratio of the mobile-liquid-like molecules. In the topmost layer of the rubber interface, the ratio of liquid-like molecules linearly increases, while that in the vapor interface shows an overall higher liquid-like ratio, yet saturating at above 259 K (Fig. 3D). In Layer 2, the rubber interface shows lower liquid-like ratios than the vapor interface below 267 K (Fig. 3E). However, this relationship switched at 269 K, with the liquid-like ratios of the rubber interface approaching those in Layer 1. This trend is consistent with the structural disordering (Fig. 2A and C). Layer 3 shows a similar trend to Layer 2, showing an unexpectedly high liquid-like ratio at the rubber interface (Fig. 3F). These results show that, at the rubber interface, molecular dynamics are restrained due to the presence of rubber molecules. However, near the melting point, the topmost layer mixes with the rubber, causing the bottom layers to locate near the interface and allowing them to behave similarly to the previous topmost layers. Despite the similarity in density profiles at 269 K between Layer 3 at the rubber interface and Layer 2 at the vapor interface (Fig. 2C), the liquid ratio was approximately 0.7 for the former and nearly 0.9 for the latter, indicating comparable but distinct molecular dynamics in the two layers.
Figure 4: Analysis of the atomistic structure of the ice–rubber interface. (A) Snapshot of Layer 1 at 269 K, showing water molecules as van der Waals spheres and rubber atoms as a surface representation. (B) Root-mean-square fluctuation (RMSF) of surface rubber atoms in Layer 1. Error bar shows standard deviation of the mean values computed using 5 ns time-blocks. (C) Representative snapshot of a region within Layer 1, where the densities of water and rubber are comparable, at 254 and 267 K. Water molecules are shown as oxygen atom positions and are colored by the score $g(\bm{x})$, which quantifies deviation from bulk ice. Higher values (yellow) indicate greater mobility. Rubber atoms are shown as green spheres.
3.3 Coupling of the dynamics of rubber molecules and premelting layers
We have shown the rubber–induced effects on ice premelting layers, which are attributed to the molecular interactions and nanoscale surface structure of the rubber. We therefore conducted a detailed analysis of the atomistic interface of the rubber. The snapshot (Fig. 4A) shows that, despite the hydrophobic nature of the rubber chains, water molecules penetrate the rubber matrix.
We hypothesized that, as the ice–rubber interface is tightly packed, the dynamics of rubber and ice molecules become coupled. To test this, we computed the RMSF of the rubber atoms within Layer 1 (Fig. 4B). The fluctuations increased steeply with temperature rise, showing a marked rise above 264 K, consistent with the temperature dependence of water dynamics at the ice–rubber interface (Fig. 3B). Furthermore, the spatial distributions of rubber and water molecules within Layer 1 indicate that the packing was looser at 267 K than at 254 K (Fig. 4C), leading to increased rubber mobility. Nevertheless, rubber atoms, mobile water molecules, and immobile water molecules each formed distinct clusters within the layer at both temperatures.
4 Discussion
Tribology of ice–rubber interface is influenced by multiple factors, including frictional heating, pressure melting, ice adhesion, and premelting layers, yet a comprehensive understanding of these effects remains elusive. Here, we aimed to clarify the rubber-induced effect on the ice premelting. From the atomistic molecular dynamics simulations, we have shown that the rubber, despite its hydrophobicity, induces structural disorder of the ice interface, enhancing the ice premelting. In contrast, based on the diffusion coefficient and liquid-like ratio, the premelting layer at the ice–rubber interface shows lower mobility compared to that at the ice–vapor interface. However, at temperatures approaching the melting point (267 to 269 K), a drastic increase in molecular dynamics was observed at the rubber-ice interface, accompanied by enhanced rubber molecular dynamics. These results demonstrate that ice and rubber mutually influence each other’s structure and dynamics, resulting in an ice-rubber interfacial layer with unique temperature-dependent properties.
Why do hydrophobic polymers enhance premelting? Previous studies have proposed a depletion mechanism 47, 48, 26, in which water molecules are excluded from the hydrophobic polymer surface due to repulsive excluded volume effects, forming a water depletion layer. In contrast, our results show that water molecules can penetrate nanoscale voids within the polymer structures, where the roughness scale is comparable to the bilayer thickness of ice. This suggests that the nanoscale morphology of the polymer, via the excluded volume effect, disrupts hydrogen-bond networks of the water molecules within the premelting layers 49. Notably, the substantial increase in rubber flexibility observed in our study is likely triggered by bilayer melting, which occurs a few kelvins below the melting point at the ice-vapor interface 14. As a result, the topmost layer becomes a mixed rubber–ice state, encouraging lower layers (Layer 2 and Layer 3) to behave as if they are the topmost interface layer.
Ice–rubber tribology is a multiscale phenomenon, and our work provides fundamental molecular-level insights into the rubber–ice interface. We speculate that the premelting layers influence both the effective friction coefficient and adhesion to ice. Bridging these molecular-level understanding with macroscopic observations may contribute to the development of functional polymer materials and their industrial applications.
Acknowledgments
This work was supported by JST, CREST (Grant No. JPMJCR2093), and in part by MEXT under the “Program for Promoting Research on the Supercomputer Fugaku” (Grant No. JPMXP1020230325). The authors are grateful to Prof. K. Kurihara (Tohoku University) and Prof. M. Mizukami (Tohoku University) for helpful discussions and valuable comments. I.Y. was supported by a Grant-in-Aid for JSPS Fellows (Grant No. JP23KJ1918).
Supporting Information
Figure S 1: Density profiles of water oxygen atom (Ice O) and rubber atoms (Rubber) at 254, 259, 264 and 269 K.
Temperature [K] Vapor Rubber
254 0.854 0.771
259 0.862 0.810
264 0.861 0.792
267 0.878 0.865
269 0.892 0.876
Table S 1: Diffusion exponent, $\alpha$ ($\rm MSD\propto t^{\alpha}$), at ice–vapor interface (Vapor) and ice–rubber interface (Rubber).